OpenStudy (anonymous):

prove that sec 8A-1 / sec 4A-1 = tan 8A/tan 2A.

OpenStudy (anonymous):

tan^x + 1 = sec^2 x

OpenStudy (anonymous):

tan^2 x + 1 = sec^2 x

OpenStudy (anonymous):

what is this ?

OpenStudy (anonymous):

it is a formula

OpenStudy (anonymous):

sec 8A = sec 2(4A)

OpenStudy (anonymous):

how do i use it ?

OpenStudy (anonymous):

aha.... got it..

OpenStudy (anonymous):

Sec 8A = 1/cos 8A

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

nw Cos 8A = Cos 2 (4A)

OpenStudy (anonymous):